OpenStudy (anonymous):

(Calculus) One side of a right triangle is known to be 20 cm long and the opposite angle is measured as 30 degrees, with a possible error of measurement of 1 degree. Use differentials to estimate the error in computing the length of the hypotenuse. No idea how to approach this problem. Missed this lesson in class...

OpenStudy (alekos):

no takers?

OpenStudy (alekos):

we know that sin(theta)= 20/H where θ = pi/6 therefore H = 20/sinθ so now we find dH/dθ

OpenStudy (alekos):

see how you go from here

OpenStudy (anonymous):

ah yes I see.. Is it as simple as that?? \[dH=\frac{20\cos \theta}{\sin^2 \theta} \times d \theta\]

OpenStudy (anonymous):

Is this correct?

OpenStudy (anonymous):

what is d theta though? I know i have to plug in 30 degrees to the other thetas.

OpenStudy (alekos):

you have to be very careful with dθ. In calculus always work in radians, so dθ is .0175 radians

OpenStudy (alekos):

.0175 radians is equivalent to 1 degree

OpenStudy (anonymous):

Oooh. I see. But if I keep it in degrees, it should come out to (I corrected my equation from before)\[dH=\frac{-20\cos (30)}{\sin^2 (30)}*\frac{\pi}{180}?\]

OpenStudy (anonymous):

well pi over 180 is basically .0175 haha. don't know if this is the right way though..

OpenStudy (alekos):

this will work but its always best to stick to radians

OpenStudy (alekos):

its correct. well done

OpenStudy (anonymous):

so instead of 30, I should use pi over 6?

OpenStudy (alekos):

strictly speaking yes

OpenStudy (anonymous):

okay awesome. It makes better sense to do radians now that I'm thinking more thoroughly about it. Thanks so much. (: This will help on my finals.

OpenStudy (alekos):

no problem. you should get an answer of dH = 1.212 cm

OpenStudy (alekos):

great problem. thanks for posting